OpenStudy (anonymous):

add, simplify if possible x+9/x-8 + x-2/8-x + 7(x+1)/x-8

OpenStudy (anonymous):

\[\frac{x+9}{x-8}+\frac{x-2}{8-x}+\frac{7(x+1)}{x-8}\]Like this? :-D

OpenStudy (anonymous):

just like that!

OpenStudy (anonymous):

\[\frac{x+9}{x-8}+\frac{x-2}{8-x}+\frac{7(x+1)}{x-8}\] \[=\frac{x+9}{x-8}-\frac{x-2}{x-8}+\frac{7(x+1)}{x-8}\] \[=\frac{x+9 -x + 2 + 7x + 7}{x-8}\] \[=\frac{7x+18}{x-8}\]